A normally distributed data has a mean of 200 and a standard deviation of 20.What are the two extreme scores, outside of which only 5% of the distribution are expected to fall?

1
Expert's answer
2022-05-06T11:19:01-0400
"P(X<x_1)=P(Z<\dfrac{x_1-200}{20})=0.025"




"\dfrac{x_1-200}{20}=-1.96"

"x_1=160.8"



"P(X>x_2)=1-P(Z\ge\dfrac{x_2-200}{20})=0.025"





"\dfrac{x_2-200}{20}=1.96"

"x_2=239.2"



160.8 and 239.2